1- a Bullet hit a wooden plate with thickness 20 cm by velocity 500 m/s , then the bullet comes out from the other side of the plate with acceleration 3*105 m/s2 .

calculate

a) the velocity of the bullet after coming out from the plate

b) the needed plate thickness to stop the bullet
Answer #1

. that a=3x10² m Answer: - Given u=500m/ d=20cm 20-20m, ca) uplu zad = (50)– (2) (3x1080.2) = 361 m V=361 m 361 2)(3x105) b)